OpenStudy (anonymous):

if the mean of this data is 95 , find the missing data 103, 99, 90, 86, 105, 96,_

OpenStudy (anonymous):

If the mean of 103, 99, 90, 86, 105, 96, Y = 95, then Y would be: (103 + 99 + 90 + 86 + 105 + 96 + Y) / 7 = 95 (579 + Y) / 7 = 95 Multiply both sides by 7; so: (579 + Y) / 7 x Y = 95 x 7 579 + Y = 665 Y = 665 ? 579 Y = 86